OpenStudy (anonymous):

8x+8/12x+60 * 4x+20/2x^2-2

OpenStudy (anonymous):

the ^ is a square root

OpenStudy (shadowfiend):

\[8x + \frac{8}{12}x + 60(4)x + \frac{20}{2}x^2 - 2\] The first thing you want to do is simplify any fractions and do any multiplications you see: \[\frac{8}{12} \rightarrow \frac{2}{3}\] \[\frac{20}{2}\rightarrow 10\] Can you rewrite the equation with those done?

OpenStudy (radar):

I think it will reduce to \[4\over3(x-1)\]

OpenStudy (anonymous):

That was right!! Thanks :)

OpenStudy (anonymous):

I have another one that I am having problems with if u have a min

OpenStudy (radar):

ok

OpenStudy (anonymous):

I have to figure out how to write them

OpenStudy (anonymous):

bare with me

OpenStudy (radar):

Yes it is important to remove any ambiguity.

OpenStudy (anonymous):

|Each figure represents a Real number and no denominators are 0. Choose the correct statement and briefly explain why the others are not correct: a) |dw:1320883621475:dw

OpenStudy (anonymous):

I tried to draw something but it didnt work

OpenStudy (anonymous):

I guess

OpenStudy (radar):

The drawing did not materialize on my pc.

OpenStudy (anonymous):

8/x+4 - 3x/3x+12 + 1/2x+8

OpenStudy (shadowfiend):

Sorry, looks like you deleted the |dw at the beginning and the | at the end. Or a glitch may have caused it.

OpenStudy (shadowfiend):

I'll let radar take this one, but you may want to consider asking it as a separate question on the left, so you can give him a medal for both answers :)

OpenStudy (anonymous):

Ok. Thanks I am new to this site.

OpenStudy (shadowfiend):

No problem! Welcome! < and I'm getting out of the way now ;) >

OpenStudy (anonymous):

I think radar left too :(

OpenStudy (radar):

The problem could be looked at different ways is this the problem: 8 3x 1 -------- - -------- + ------- x+4 3x+12 2x + 8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry not sure how to get them to look like that

OpenStudy (radar):

If it is note that the denominators have in common x+4. Ok rewrite the denominators as follows: 8 3x 1 ---- - ------ + ------- x+4 3(x+4) 2(x+4)

OpenStudy (radar):

The 2nd term can be simplified by dividing the numerator by 3 getting: x ---------- x+4

OpenStudy (radar):

Subtract the first two terms giving: 8-x ---- x+4 now add the last term giving 8.5-x ------ x+4 That is 8 1/2 - x -------- x+4

OpenStudy (radar):

That looks crazy I may have erred.

OpenStudy (anonymous):

Yeah it cant be because we arent doing anything that would give us 8.5

OpenStudy (anonymous):

it should be whole numbers, but thanks for trying. These are HARD!

OpenStudy (anonymous):

I have been working on this all day :(

OpenStudy (radar):

@shadowfiend where did I go wrong?

OpenStudy (radar):

I think you have to use the denominator 2(x+4) to do the addition. I got 9-x ---- x+4

OpenStudy (anonymous):

it says its wrong